OpenStudy (anonymous):

Question 1.1. Consider the system of equations: 3x + 2y = 18 and x – 2y = 14. What property is used to transform the equation x – 2y = 14 into – 3x + 6y = – 42? (Points : 1) addition property of equality substitution property of equality multiplication property of equality division property of equality

Mehek (mehek14):

x*3 = 3x 2y*3 = 6y 14 *3 = 42 can you sort of see what the answer should be?

OpenStudy (anonymous):

sorry but no

Mehek (mehek14):

what do you do each time? you multiply by three what property is it then?

OpenStudy (anonymous):

Multiplication property of equality?

Mehek (mehek14):

yes

OpenStudy (jdoe0001):

hehe "Multiplication property of equality" <---- what does that mean?

OpenStudy (anonymous):

ok Thank u so much

Mehek (mehek14):

when you multiply both sides of the equation by the same number @jdoe0001

OpenStudy (jdoe0001):

well @Mehek14 yes, but he/she's just picking a choice, rather than understand what's being done

OpenStudy (jdoe0001):

picking prefabricated options and actually understanding, is not exactly the same thing

Mehek (mehek14):

true